The Amtrak auto train travels nonstop from lorton, Virginia to sanford, Florida. June wishes to bring two cars down to Florida for the winter, so she decides to send one by auto train and drive the other one down. The train travels 600 kilometers in the time it takes June to drive 400 kilometers. If the average speed of the train is 40 kilometers faster than the speed of junes car, find the speeds of the train and car.
Answer by lwsshak3(11628)

The Amtrak auto train travels nonstop from lorton, Virginia to sanford, Florida. June wishes to bring two cars down to Florida for the winter, so she decides to send one by auto train and drive the other one down. The train travels 600 kilometers in the time it takes June to drive 400 kilometers. If the average speed of the train is 40 kilometers faster than the speed of junes car, find the speeds of the train and car.
..
let x=speed of car (km/hr)
then x+40 = speed of train (km/hr)
Time of travel =distance/speed (same for car & train)
400/x=600/x+40
600x=400x+16000
200x=16000
x=80 km/hr
ans:
speed of car=80 km/hr
speed of train=80+40=120 km/hr